OpenStudy (anonymous):

Help me if possible!! Solve for x: 25x2 - 36 = 0 (2 points) x = 6, 5 x = -6, -5 x = , x = - x = , x = -

pooja195 (pooja195):

What are the last 2 choices?

OpenStudy (anonymous):

for c it is 36 over 25 and negative 36 over 25 (as in the whole equation. for d it is 6 over 5 and negative 6 over 5 (whole equation is neg)

OpenStudy (anonymous):

sorry it idnt come through

pooja195 (pooja195):

Add 36 to both sides. \[\huge~\rm~25x^2−36+36=0+36\]\[\huge~\rm~25x^2=36\]

pooja195 (pooja195):

Do you know what to do from here?

OpenStudy (anonymous):

do i divide??

OpenStudy (anonymous):

wait no

pooja195 (pooja195):

yes! :D I added 36 to both sides because we are trying to ISOLATE x :) Next step is indeed to divide!

OpenStudy (anonymous):

oksy

OpenStudy (anonymous):

is the answer d?

pooja195 (pooja195):

Thats correct

OpenStudy (anonymous):

thank you so much.

pooja195 (pooja195):

You simplyfy the orginal answer which is 36/25

pooja195 (pooja195):

No problem :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

\[25x^2 - 36 = 0\\(5x)^2-6^2=0\\(5x+6)(5x-6)=0\ [use\ a^2-b^2=(a+b+(a-b)]\\ \implies\ either\ (5x+6)=0\ or\ (5x-6)=0\\ \implies\ either\ 5x=-6\ or\ 5x=6\\ \implies\ either\ x=\frac{ -6 }{ 5 }\ or\ x=\frac{ 6 }{5 }\]